According to a study done last year, the average monthly expenses for mobile phone loads of college students in San Mateo, Rizal was P 400.00. A statistics student believes that this amount has decreased since January of this year. Is there a reason to believe that this amount has decreased if a random sample of 50 students has an average monthly expense for mobile phone loads of P 380.00? Use .05 level of significance. Assume the population standard deviation is P 75.00.  

Step 1: State the null and alternative hypotheses.

(1.) Null Hypothesis (Ho):

(2.) Alternative Hypothesis(Ha):

(3.) Write Ho in words:

(4.) Write Ha in words:

Step 2: Level of significance and the Type of test.

(5.) Level of Significance: α =

(6.) Two-tailed Test or One-tailed Test

Step 3: Compute the test statistic.

(7.) = (8.)  = ______ (9.) ẟ = ______ (10.) n = _or   s =

ComputationEncircle the formula to be used

4: Determine the critical value.


1
Expert's answer
2021-03-25T14:04:30-0400

We have that:

"\mu=400"

"\sigma=75"

"n=50"

"\bar x=380"

"\alpha=0.05"


"H_0:\mu=400"

"H_a:\mu<400"

H0 : "the average monthly expenses for mobile phone loads of college students in San Mateo, Rizal is P 400.00"

Ha : "the average monthly expenses for mobile phone loads decreased"


The hypothesis test is left-tailed.

The population standard deviation is known and the sample size is large (n>30) so we use z-test.

The critical value for α = 0.05 is Z0.05 = –1.64

The critical region is z < –1.645.

Test statistic:


"Z_{test}=\frac{\bar x - \mu}{\frac{\sigma}{\sqrt n}}=\frac{380- 400}{\frac{75}{\sqrt 50}}=-1.89"


Since –1.89 < –1.645 thus the Ztest falls in the rejection region we reject the null hypothesis.

At the 5% significance level the data do provide sufficient evidence to support the claim. We are 95% confident to conclude that the average monthly expenses for mobile phone loads decreased.